Problem 2.7 Part A If a pitcher throws a 86 mph fastball, how much time elapses from when the ball leaves the pitchers hand until the ball reaches the batter? feet from the batter Express your answer to two significant figures and include the appropriate units. My Answers Give Up Provide Feedback Continue
Answer #1

1 feet = 0.3048 m

thus 60 feet = 60* 0.3048 = 18.288 m

1 mile = 1609.344 m

thus 86mph = 86* 1609.344 / 3600 = 38.445 m/s

now time = distance / speed = 18.288/38.445 = 0.476 seconds